OpenStudy (loser66):

Help me, please. I confused f:R\0 --> R , f(x) = 1/x^2 E =[1,2] find f(E) and f^- (E)

OpenStudy (helder_edwin):

if \[\large 1\leq x\leq2 \] then \[\large 1\leq x^2\leq4 \] \[\large 1\geq\frac{1}{x^2}\geq\frac{1}{4} \] so \[\large f(E)=[1/4;1] \]

OpenStudy (helder_edwin):

I see.

OpenStudy (helder_edwin):

OK. it makes no sense to me, but i understand. U have a typo: \[\large g^-(E)=[-2;-1]\cup[1;2] \]

OpenStudy (helder_edwin):

yes. so u would get \[\large f^-(E)=[-1;-1/\sqrt{2}]\cup[1/\sqrt{2};1] \]

OpenStudy (helder_edwin):

\[\large g^-(E)=[-\sqrt{2};-1]\cup[1;\sqrt{2}] \] \[\large h^-(g^-(E))=h^-([-\sqrt{2};-1]\cup[1;\sqrt{2}])= \] \[\large =h^-([-\sqrt{2};-1])\cup h^-([1;\sqrt{2}]) \] \[\large =[-1;1/\sqrt{2}]\cup[1/\sqrt{2};1] \]

OpenStudy (helder_edwin):

sorry. but i have to go.

OpenStudy (loser66):

let g(x) = 1/x , so if E = \(\{ 1\leq x\leq 2\}\), then g(E) = \(\{ 1/2\leq x\leq 1\}=[1/2,1]\) \(g^-(x) = 1/x \) so \(g^-(E) = [1/2,1]\) am I right?

OpenStudy (kirbykirby):

I feel bad I stuck around here for awhile lol. But what you wrote seems right to me so far. I'm not sure how helder defined g(x) in their response? is it the same g(x) you used

OpenStudy (loser66):

Pleaaaaaaaaaaaase, check,

OpenStudy (loser66):

:) @kirbykirby I got mad myself when I can't solve it by myself.

OpenStudy (kirbykirby):

. The way I did it was the following, and gave the same result as @helder_edwin \(1\le x \le 2\), Let \(\dfrac{1}{x^2}=f(x) \implies\pm \sqrt{\dfrac{1}{x}}=f^{-1}(x)\) \(\implies \dfrac{1}{\sqrt{x}}\) OR \( -\dfrac{1}{\sqrt{x}}\) so, consider the positive branch first ... transform your interval into the given format. \(1\le x \le 2 \implies 1\ge \dfrac{1}{x}\ge \dfrac{1}{2} \\ ~ \\ \implies 1 \ge \dfrac{1}{\sqrt{x}} \ge \dfrac{1}{\sqrt{2}} \\ \implies \dfrac{1}{\sqrt{2}}\ \le \dfrac{1}{\sqrt{x}} \le 1\) Now the negative branch: It's the same process, but at the last line: \(-\dfrac{1}{\sqrt{2}}\ \ge -\dfrac{1}{\sqrt{x}} -\ge 1\) \( \implies -1 \le -\dfrac{1}{\sqrt{x}} \le -\dfrac{1}{\sqrt{2}}\) So overall: \(\left[-1,-\dfrac{1}{\sqrt{2}}\right] \cup \left[\dfrac{1}{\sqrt{2}},1 \right]\)

OpenStudy (kirbykirby):

@Loser66

OpenStudy (kirbykirby):

(your attachment wouldn't load for some reason o.O)

OpenStudy (loser66):

And it is the same answer with me, yeahhhhhhhhhhhhhh. Thank you

OpenStudy (kirbykirby):

ah good :) awesome